# vec2wav 2.0: Advancing Voice Conversion via Discrete Token Vocoders thanks:

### v2

@@ II-B Adaptive Snake Activation
where $T$ is a linear transform followed by $\tanh$ activation, and operations in (2) are all element-wise. $T(\bm{s})$ is discounted by $1/2$ on the magnitude part for numerical stability. To save parameters, we apply the same $T$ transformation to both magnitude and frequency. In this way, the learnable parameter for each adaptive Snake is $\theta=\{\bm{\alpha},\bm{\beta},W,\bm{b}\}$, and the target timbre information can be effectively injected in every layer of the vocoder via adaptive activations, with strengthens the timbre controllability to a considerable extent.
@@ II-C Content and Prompt Features
Content Features   We use the off-the-shelf vq-wav2vec [8] SSL model for extracting the discrete content representation to be re-synthesized. The discrete tokens are extracted from the quantizer output before the feature aggregator, which is a two-group integer index sequence. We favor this representation because a lot of speaker timbre information is removed due to the information bottleneck, while most of the phonetic pronunciation and prosody are retained [33]. Also, compared to HuBERT [9]-style Transformer SSL models, vq-wav2vec is free of manual clustering and is also fully convolutional with a certain receptive field. This produces a representation that is unaware of the total sequence length, keeping consistent results for a given window. This consistency also shows potential for cross-lingual conversion, as its language-agnostic property has been successfully applied in multilingual TTS [34]. Although there exists measurable speaker timbre leakage in the discrete tokens [35, 36, 33], the vec2wav 2.0 vocoder exhibits strong timbre controllability, so that competitive VC can still be achieved.
@@ IV Conclusion
We present a novel VC method, vec2wav 2.0, based on the re-synthesis of speech discrete tokens. It takes advantage of SSL features in both content and timbre representations and enhances CTX-vec2wav in architectural designs. The adaptive Snake activation technique is proposed to better incorporate timbre into waveform reconstruction. The resulting model achieves remarkable performance on intra and cross-lingual VC tasks. We believe vec2wav 2.0 has promising impacts on the future LLM-based speech generation paradigm. Future efforts are needed in improving the intelligibiltiy and prosody preservation of the proposed method.

### v3

@@ II-B Adaptive Snake Activation
where $T$ is a linear transform followed by $\tanh$ activation, and operations in (2) are all element-wise. $T(\bm{s})$ is discounted by $1/2$ on the magnitude part for numerical stability. To save parameters, we apply the same $T$ transformation to both magnitude and frequency. In this way, the learnable parameter for each adaptive Snake is $\theta=\{\bm{\alpha},\bm{\beta},W,\bm{b}\}$, and the target timbre information can be effectively injected in every layer of the vocoder via adaptive activations, which strengthens the timbre controllability to a considerable extent.
@@ II-C Content and Prompt Features
Content Features   We use the off-the-shelf vq-wav2vec [8] SSL model for extracting the discrete content representation to be re-synthesized. The discrete tokens are extracted from the quantizer output before the feature aggregator, which is a two-group integer index sequence. We favor this representation because a lot of speaker timbre information is removed due to the contrastive criterion, while most of the phonetic pronunciation and prosody are retained [33]. Also, compared to HuBERT [9]-style Transformer SSL models, vq-wav2vec is free of manual clustering and is also fully convolutional with a certain receptive field. This produces a representation that is unaware of the total sequence length, keeping consistent results for a given window. This consistency also shows potential for cross-lingual conversion, as its language-agnostic property has been successfully applied in multilingual TTS [34]. Although there exists measurable speaker timbre leakage in the discrete tokens [35, 36, 33], the vec2wav 2.0 vocoder exhibits strong timbre controllability, so that competitive VC can still be achieved.
@@ IV Conclusion
We present a novel VC method, vec2wav 2.0, based on the re-synthesis of speech discrete tokens. It takes advantage of SSL features in both content and timbre representations and enhances CTX-vec2wav in architectural designs. The adaptive Snake activation technique is proposed to better incorporate timbre into waveform reconstruction. The resulting model achieves remarkable performance on intra and cross-lingual VC tasks. We believe vec2wav 2.0 has promising impacts on the future LLM-based speech generation paradigm. Future efforts are needed in improving the intelligibility and prosody preservation of the proposed method.